# Ratio asymptotics and zero density for orthogonal polynomials with varying Verblunsky coefficients

### v1

@@ 6. Examples
The Rogers–Szegő polynomials [simon_opuc1, Ex. 1.6.5] correspond to the choice of Verblunsky parameters $\alpha_{n}=(-1)^{n}\gamma^{n+1}$ with $\gamma\in(0,1)$; see also [ismail_09, Chap. 17]. The orthogonality measure is then given by
$\frac{{\rm d}\mu}{{\rm d}\theta}(e^{{\rm i}\theta})=\frac{1}{\sqrt{2\pi a}}\sum_{j=-\infty}^{\infty}e^{-(\theta-2\pi j)^{2}/(2a)},$ (6.7)
@@
Instead of analyzing the case $\alpha_{n,N}=(-1)^{n}\zeta^{(n+1)/N}$, which corresponds to the choice $\gamma_{n}=\zeta^{1/N}$, we consider the rotated case $\alpha_{n,N}=\zeta^{(n+1)/N}$. Let us therefore study the latter case. As $n/N\to t$, we obtain the setting of Theorem 18 and Proposition 23 with strictly decreasing $f(s)=\zeta^{s}$.
@@
$\mathop{\mathrm{supp}}\nolimits\sigma_{t}=\{e^{{\rm i}\theta}\mid\theta\in[2\arcsin\zeta^{t},2\pi-2\arcsin\zeta^{t}]\}$
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\!\left(e^{{\rm i}\theta}\right)=\frac{1}{2\pi t}\int_{\log_{\zeta}\sin(\theta/2)}^{t}\frac{\sin\left(\theta/2\right)}{\sqrt{\sin^{2}\left(\theta/2\right)-\zeta^{2s}}}{\rm d}s=-\frac{1}{2\pi t\log\zeta}\int_{\zeta^{t}/\sin(\theta/2)}^{1}\frac{{\rm d}u}{u\sqrt{1-u^{2}}},$
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)=\frac{1}{2\pi}\left(1-\frac{1}{t}\log_{\zeta}\left(\sin(\theta/2)+\sqrt{\sin^{2}(\theta/2)-\zeta^{2t}}\right)\right),$ (6.8)
@@
The limiting zero density for the varying Rogers–Szegő polynomials with $\alpha_{n,N}=(-1)^{n}\zeta^{(n+1)/N}$ is then $\sigma_{t}(e^{{\rm i}(\theta-\pi)})$ where $\sigma_{t}$ is  (6.8).
Figure 2. Density (6.8) plotted for $q\in\{0.1,0.3,0.5,0.7,0.9\}$, $q:=\zeta^{2}$.
@@
Let $f(s)=\sqrt{1-s^{2}}$ and $t\leq 1$. Then $f$ is strictly decreasing and a straightforward computation of the integral in (6.2) results in the formula
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)=\frac{1}{2\pi t}\sin\left(\theta/2\right)\left[\log\left(t+\sqrt{t^{2}-\cos^{2}\left(\theta/2\right)}\right)-\log\left|\cos\left(\theta/2\right)\right|\right],$ (6.9)
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)$ $=\frac{1}{2\pi t}\int_{0}^{t\theta/\pi}\frac{\sin\left(\theta/2\right)}{\sqrt{\sin^{2}\left(\theta/2\right)-\sin^{2}\left(\pi s/2t\right)}}{\rm d}s$
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)=\frac{1}{\pi^{2}}\sin\left(\theta/2\right)K\left(\sin\left(\theta/2\right)\right),$ (6.11)

### v2

@@ 6. Examples
The Rogers–Szegő polynomials [simon_opuc1, Ex. 1.6.5] correspond to the choice of Verblunsky parameters $\alpha_{n}=(-1)^{n}\gamma^{n+1}$ with $\gamma\in(0,1)$; see also [ismail_09, Chap. 17]. Their orthogonality measure is supported on $\partial\mathbb{D}$} with the density
$\frac{{\rm d}\mu}{{\rm d}\theta}({\rm e}^{{\rm i}\theta})=\frac{1}{\sqrt{2\pi a}}\sum_{j=-\infty}^{\infty}{\rm e}^{-(\theta-2\pi j)^{2}/(2a)},$ (6.5)
@@
Instead of analyzing the case $\alpha_{n,N}=(-1)^{n}\zeta^{(n+1)/N}$, which corresponds to the choice $\gamma_{n}=\zeta^{1/N}$, we consider the rotated case $\alpha_{n,N}=\zeta^{(n+1)/N}$. Let us therefore study the latter case. As $n/N\to t$, we obtain the setting of Theorem 18 and Proposition 23 with strictly decreasing function $f(s)=\zeta^{s}$.
@@
$\mathop{\mathrm{supp}}\nolimits\sigma_{t}=\{{\rm e}^{{\rm i}\theta}\mid\theta\in[2\arcsin\zeta^{t},2\pi-2\arcsin\zeta^{t}]\}$
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\!\left({\rm e}^{{\rm i}\theta}\right)=\frac{1}{2\pi t}\int_{\log_{\zeta}\sin(\theta/2)}^{t}\frac{\sin\left(\theta/2\right)}{\sqrt{\sin^{2}\left(\theta/2\right)-\zeta^{2s}}}{\rm d}s=-\frac{1}{2\pi t\log\zeta}\int_{\frac{\zeta^{t}}{\sin(\theta/2)}}^{1}\frac{{\rm d}u}{u\sqrt{1-u^{2}}},$
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left({\rm e}^{{\rm i}\theta}\right)=\frac{1}{2\pi}\left[1-\frac{1}{t}\log_{\zeta}\!\left(\sin(\theta/2)+\sqrt{\sin^{2}(\theta/2)-\zeta^{2t}}\right)\right],$ (6.6)
@@
The limiting zero density for the varying Rogers–Szegő polynomials with $\alpha_{n,N}=(-1)^{n}\zeta^{(n+1)/N}$ is then $\sigma_{t}({\rm e}^{{\rm i}(\theta-\pi)})$, where $\sigma_{t}$ is  (6.6).
Figure 2. Density (6.6) plotted for $t=1$ and $\zeta\in\{0.1,0.3,0.5,0.7,0.9\}$.
@@
Let $f(s)=\sqrt{1-s^{2}}$ and $0<t\leq 1$. Then $f$ is strictly decreasing and a straightforward computation of the integral in (6.2) results in the formula
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left({\rm e}^{{\rm i}\theta}\right)=\frac{\sin\left(\theta/2\right)}{2\pi t}\left[\log\left(t+\sqrt{t^{2}-\cos^{2}\left(\theta/2\right)}\right)-\log\left|\cos\left(\theta/2\right)\right|\right],$ (6.7)
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}({\rm e}^{{\rm i}\theta})$ $=\frac{1}{2\pi t}\int_{0}^{t\theta/\pi}\frac{\sin\left(\theta/2\right)}{\sqrt{\sin^{2}\left(\theta/2\right)-\sin^{2}\left(\pi s/2t\right)}}{\rm d}s$
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}({\rm e}^{{\rm i}\theta})=\frac{\sin\left(\theta/2\right)}{\pi^{2}}\,K\!\left(\sin\left(\frac{\theta}{2}\right)\right),$ (6.9)